# Assessing the impact of variance heterogeneity and misspecification in mixed-effects location-scale models

### v1

@@ 1 Background
Longitudinal studies are critical to answering research questions in many fields because of their ability to highlight trends and changes over time. In medicine, these analyses are fundamental to capturing a disease’s evolution through different stages of life, as well as measuring the changes in treatments’ effects. In social sciences, they are used to track the evolution of socioeconomic phenomena, such as the impact of where one lives on economic outcomes [22], and the associated impact of policies on these outcomes.
@@
In the analysis of longitudinal data, linear mixed models (LMMs) emerged as a popular option to model repeated measures for the same statistical unit, such as individuals or countries, because of their interpretability and the availability of implementations in multiple statistical software [18, 3]. The key advantage of LMMs is that one can model the population average as a function of time and covariates – these population-level relations are known as fixed effects –, and individual-specific predictions, known as random effects, which capture the correlation between the measures for the same unit and quantify the deviation with respect to the population average. Random effects are assumed to be drawn from a common distribution (usually normal) with zero mean.
@@
An underlying assumption of LMMs is homoscedasticity: the residual variance is assumed constant across all observations and covariates. In other words, the effect of unobserved covariates, which is captured by the error term, contributes to variation in equal ways for all units. While some options to validate this assumption exist (e.g.L̃evene’s test [8] for homogeneity of variance, or fitted-residual plots to graphically assess the absence of unexplained patterns), these are not commonly reported in practice. While heteroscedasticity does not necessarily bias fixed effect estimates, ignoring this heterogeneity can impact variance estimates and, consequently, result in Type I and II errors [13, 14].
Modelling heterogeneity may even reflect important population and individual characteristics. For instance, in cardiovascular research, the notion of blood pressure variability [15] is becoming more commonly used as a risk factor for stroke and other events, as large fluctuations over time might be indicative of arterial health and other issues. This could lead to finding new treatments which do not only bring blood pressure to a normal range but also stabilise it around some average values.
@@
In the statistics literature, multiple models model the variance as well as the mean. Hedeker et al [4, 5] introduced a mixed-effect location-scale model (MELSM) to model both the location parameter (the mean) and the scale parameter (the residual standard deviation) as a function of covariates and random effects. The model was initially applied to quantify the within-subject heterogeneity in ecological momentary assessments to study mood variation [11], but then has been applied to other domains, e.g. cystic fibrosis studies [9]. A general modelling framework is provided in gamlss [12] and its Bayesian version (bamlss, 17) with a large range of distributions, all parameters can be modelled using fixed and random effects (although not many examples explicitly model within-individual variability using random effects). All these options fall under the umbrella of distributional regression models, where the conditional distribution of the outcome given the covariates is studied via its parameters.
@@
Despite the existence of such solutions, modelling heteroscedasticity is not often done in practice. Walters et al [20] discussed how software availability and a lack of best practices in specifying the scale model may explain this lack of adoption. With recent user-friendly software developments such as [2] and [12], only the second barrier remains. Further, methodological challenges exist when using MELSMs. For example, although the problem of power analysis and testing the homogeneity of variance in MELSM was the object of some recent studies [20, 21], the relationship between location and scale is not fully understood, and how the misspecification of one affects the other.
@@
The paper is structured as follows. Section 2 introduces our notations and the studied MELSM. Section 3 describes the simulation setting. In Section 4, we study six common practices in specifying the scale model and explore how misspecifications affect MELSM’s estimates. Finally, in Section 5, we discuss the insights from the simulations, their limitations, and potential further developments.
@@ 2 Mixed-effect location-scale model
Consider the normally-distributed outcome $y_{i,j}\in\mathbb{R}$ for subject $i\in[1,N]$ at encounter $j\in[1,n_{i}]$. A linear mixed effect model assumes the following relation between outcome and covariates:
$y_{i,j}=x_{i,j}^{yT}\beta^{y}+z_{i,j}^{yT}u^{y}_{i}+\epsilon_{i,j}$
A linear mixed-effect model assumes that $y_{i,j}\sim\mathcal{N}(x_{i,j}^{yT}\beta^{y}+z_{i,j}^{yT}u^{y}_{i},\omega^{2})$, where the standard deviation $\omega$ is constant for all observations (homoscedasticity). The MELSM relaxes this assumption by specifying a model for the scale, which can depend on random effects, as well as covariates, i.e.:
$\omega_{i,j}:=\exp(x_{i,j}^{\omega T}\beta^{\omega}+z_{i,j}^{\omega T}u^{\omega}_{i})$
where $\beta^{\omega}$ gives some insights on the contribution of the different covariates on the variance. We model the logarithm of the standard deviation to ensure non-negativity. The random effect vectors $u^{y}_{i}$ and $u^{\omega}_{i}$ are to be interpreted as subject-specific deviations from the population mean and standard deviation, respectively. For ease of notation, random intercepts $u^{y}_{0,i}$ and $u^{\omega}_{0,i}$ are implicitly included by setting the first column of the corresponding covariate matrices to a vector of 1s, while random slopes $u^{y}_{1,i}$ and $u^{\omega}_{1,i}$ are obtained by setting the column to the vector of observation times. The random effects for the location and scale models are assumed to be normally distributed with mean zero, and a correlation between random effects can be estimated. For example, in the simple setting with just a random intercept for both the location and scale models, we can specify the correlation $\rho$ to capture the relationship between the two models:
$\begin{pmatrix}u_{i}^{y}\\ u_{i}^{w}\end{pmatrix}\sim\mathcal{N}\begin{bmatrix}\begin{pmatrix}0\\ 0\end{pmatrix}\!\!,\begin{pmatrix}\sigma^{y^{2}}&\rho\sigma^{y}\sigma^{\omega}\\ \rho\sigma^{y}\sigma^{\omega}&\sigma^{\omega^{2}}\end{pmatrix}\end{bmatrix}$ (1)
@@ 3.1 Data generation
We generate simulation data based on the Primary Biliary Cholangitis (PBC) dataset, publicly available in the R package survival, commonly used as a test for longitudinal modelling. The dataset results from a clinical trial at Mayo Clinic studying the impact of D-penicillamine on PBC, an auto-immune condition affecting the liver. From the dataset, we extract four covariates at baseline: age, albumin, triglycerides (trig), and platelet.
Using these covariates, we generate a synthetic population of $N$ individuals by drawing from a normal distribution with the covariance defined by the previously extracted and standardised covariates at baseline. Note that $N$ is a parameter of the simulation and not the observed PBC population size, as one of our aims is to study how an increasing number of patients impacts our conclusions. Formally, we draw $x_{i,0}$ – where $i$ describes the index associate with the patient and 00, the baseline encounter – from a multivariate normal distribution $\text{MVN}(0,\Sigma)$ with $\Sigma$, the covariance matrix described in Table 1.
@@
where $M$ is a parameter of the simulation study representing the average number of encounters per subject across the population. For each encounter, we compute the age at this encounter as follows:
@@
and we use this covariate as the only time-varying variable while all other covariates remain constant over time. Unless otherwise specified, we consider a total of $N=200$ patients with an average number of observations $M=15$. Further, we simulate random intercepts for all individuals defined as in Equation 1. We set $\rho=0$, $\sigma^{y}=2$ and $\sigma^{\omega}=1$.
@@ 3.2 Estimands of interest
Our simulations explore the recovery of the fixed and random effects for both scale and location models. To this end, we report the box plots and the associated coverage (for 95% credible intervals) for the estimated coefficients of interest across 100 repetitions of the simulation described in Section 3.1. Formally, the coverage of the parameter $\theta$ (in percent) is defined as follows:
@@
in which $L$ and $U$ are the lower and upper bounds of the estimated $\hat{\theta}$ at the confidence level $\alpha$.
@@ 3.3 Software
All experiments are implemented in R with code made available on GitHub [^1] https://github.com/Jeanselme/IndividualVariability. We fit the MELSM models using the brms package [2], with default priors, two chains, and a burn-in of 1000 iterations for a total of 2000 iterations.
@@ 4 Results
This section describes common practices when using mixed effect models and their impact on parameters’ estimates. For each model, we report the formula in pseudo-R code for reproducibility.
@@ 4.1 Practice 1: Ignoring heteroscedasticity
By relying on LMM, practitioners typically make an assumption of homoscedasticity, ignoring potential subject-specific differences in variance. In this first setting, we compare an LMM, assuming a constant scale, with the properly specified MELSM model. The two models share the same (correct) specification for the location, with only the scale differing:
@@
In this first set of simulations, we explore the impact of varying the number of patients and observations per patients, i.e. $N\in\{100,300,500,1000\}$ and $M\in\{5,10,20\}$. These values were chosen to approximate different data amounts and frequencies as common in real longitudinal data settings. Figure 1 presents the estimated fixed effects associated with age in the location and scale models under varying numbers of individuals $N$ across 100 simulations. Similarly, Figure 2 presents the same results under varying average numbers of observations per patients. In each figure, the red vertical line indicates the true estimand’s value.
@@
The coverage in the LMM is well below the nominal level of 95%, indicating that the standard errors are underestimated. Consequently, the misspecification of the scale has a direct impact on the location.
Larger number of individuals and encounters reduce error.
Both models benefit from a larger number of individuals and encounters as shown by reduced variance in the location estimates, and reduced variance for MELSM scale estimates.
@@ 4.2 Practice 2: Misspecifying location and scale in MELSM
Figure 3: Estimated age effect for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
Figure 4: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel). For each model, coverage is reported as percentage next to the corresponding boxplot.
Figure 3 presents the fixed effect associated with the age covariate, while Figure 4 describes the estimated standard deviation of the random intercept across the simulations. On the right, we display the coverage in percent over the different simulations. For consistency, we display all method specifications in these figures. Note that some do not estimate all quantities, hence the absence of some estimates.
@@
In all figures, the correctly specified model (Correct) and the one considering all covariates (All) result in similar estimates. This shows that, with enough data, all models that include the correct covariates as a special case will lead to a correct inference. Note that, however, considering all covariates results in an increased number of parameters and associated computational time — the method considering all covariates is more than 50% slower than the correctly specified model.
@@
The age effect in the location model is well estimated across all models, although the variance of the estimates increases as the scale model departs from the true one (as observed with LMM). The largest variance is observed for No $u^{\omega}$ + Mis. $\omega$, where both a relevant covariate and the random intercept are missing in the scale model. The same model also shows, on average, lower standard deviations for the scale random intercept than the true value.
@@
In the misspecified location model Mis. $y$, the scale fixed effect estimate and random intercept are biased, despite presenting the right scale specification. As the location parameterisation does not capture the full variation of the data, the residual variance is different from the true value, and therefore, all scale estimates are biased (Appendix A provides a theoretical intuition of why this occurs).
@@
The estimated standard deviation of location random intercept can incur bias when the scale model is incorrectly specified (and therefore, the same happens in LMM when there is heteroscedasticity between the individuals). The standard deviation for the random effect of the scale model is biased in all considered scenarios unless the fixed effect structure for both location and scale are correctly specified.

### v2

@@ 1 Background
Longitudinal studies are critical to answering research questions in many fields because of their ability to highlight trends and changes over time. In medicine, these analyses are fundamental to capturing the evolution of a disease through different stages of life, as well as measuring the changes in treatments’ effects. In social sciences, they are used to track the evolution of socioeconomic phenomena, such as the impact of where one lives on economic outcomes [wilson2012truly], and the associated impact of policies on these outcomes.
@@
In the analysis of longitudinal data, linear mixed models (LMMs) emerged as a popular option to model repeated measures for the same statistical unit, such as individuals or countries, because of their interpretability and the availability of implementations in multiple statistical software [verbeke1997linear, diggle2024longitudinal]. The key advantage of LMMs is that one can model the population trajectory (i.e. the overall longitudinal trend) as a function of time and covariates. These coefficients, which apply to all statistical units, are known as fixed effects, while individual-specific coefficients, known as random effects, can capture the correlation between the observed measurements for the same unit and quantify the deviation with respect to the population trajectory. Random effects are assumed to be drawn from a common distribution (usually normal) with zero mean.
@@
A common assumption made in the LMMs literature is homoscedasticity: the residual variance is assumed constant across all observations and covariates. In other words, the effect of unobserved covariates, which is captured by the error term, contributes to variation in equal ways for all units. While some options to validate this assumption exist (e.g. Levene’s test [levene1960robust] for homogeneity of variance, or fitted-residual plots to graphically assess the absence of unexplained patterns), these are not commonly reported in practice. While heteroscedasticity does not necessarily bias fixed effect estimates, ignoring this heterogeneity can impact variance estimates and, consequently, result in Type I and II errors [rosopa2013managing, rosopa2019conditional].
Modelling variance heterogeneity at the individual level is becoming increasingly of interest in many clinical settings, as a way to quantify important population and individual characteristics in addition to average trends over time. For instance, in cardiovascular research, the notion of blood pressure variability [rothwell2010prognostic, parati2005blood] is becoming more commonly used as a risk factor for stroke and other events, as large fluctuations over time might be indicative of arterial health and other issues. Recent work has been published stating the importance of such concepts in cardiovascular studies [stevens2019utility]. This could lead to the discovery of new treatments that not only bring blood pressure into a normal range but also stabilise its variability. In some instances [muntner2015visit, yano2017visit], summary statistics, such as the standard deviation or the coefficient of variation, computed for each individual have been used as measures of variability. barrett2019estimating showed that these summary statistics are affected by measurement error and lack precision, especially in lower-sample-size settings, common to many longitudinal studies. This observation evidences that summary statistics are not effective in modelling variance heterogeneity, therefore, more sophisticated statistical models are needed to capture this heterogeneity.
@@
The statistics literature offers model-based solutions for addressing the quantification of the individual variance. hedeker2008application, hedeker2012modeling introduced a mixed-effect location-scale model (MELSM) to model both the location parameter (the mean) and the scale parameter (the residual standard deviation) as a function of covariates and random effects. MELSM effectively incorporates variance heterogeneity in the LMM framework, while reducing the reliance on the assumption of homoscedasticity (effectively, as MELSM includes LMM as a special case, the model can be used to understand whether the assumption holds). The model was initially applied to quantify the within-subject heterogeneity in ecological momentary assessments of mood variation [rast2012modeling], but has since been applied to other domains such as cystic fibrosis [palma2024demographic] and, outside biomedical applications, student learning [leckie2024mixed]. In terms of computing options, MELSM can be fitted in MLwiN [rasbash2000user] or on Mplus, by specifying the location-scale model as a multilevel structural equation model [mcneish2021specifying]. A more general modelling framework is provided in gamlss [gamlss] and its Bayesian counterpart (bamlss, umlauf2018bamlss): a large range of outcome distributions is available, and all parameters of the distribution can be modelled using fixed and random effects (although not many examples explicitly model within-individual variability using random effects). These options fall under the umbrella of distributional regression models [klein2024distributional].
@@
Despite the existence of such solutions, modelling heteroscedasticity is not often done in practice. walters2018power discussed how software availability and a lack of best practices in specifying the scale model may explain this lack of adoption. With recent user-friendly software developments such as [burkner2017brms] and [gamlss], only the second barrier remains. Further, methodological challenges exist when using MELSMs. For example, although the problem of power analysis and testing the homogeneity of variance in MELSM was the object of some recent studies [walters2018power, williams2021reliability], the relationship between location and scale is not fully understood, and how the misspecification of one affects the other.
@@
The paper is structured as follows. Section 2 introduces our notations and the studied MELSM. Section 3 describes the simulation setting. In Section 4, we study six common practices in specifying the scale model and explore how misspecifications impact MELSM’s estimates. Finally, in Section 5, we discuss the insights from the simulations, their limitations, and potential further developments.
@@ 2 Mixed-effect location-scale model
Consider the normally-distributed outcome $y_{i,j}\in\mathbb{R}$ for subject $i\in[1,N]$ observed at encounter $j\in[1,n_{i}]$ sampled from a non-informative encounter process [^1] When this assumption does not hold, one needs to model the encounter process as proposed in [gasparini2020mixed]., where $[a,b]$ denotes the interval of integers from $a$ to $b$, inclusive. The value $N$ is the number of subjects, and the value $n_{i}$ is the number of observations for subject $i$. A linear mixed effect model assumes the following relation between outcome and covariates:
$y_{i,j}=x_{i,j}^{yT}\beta^{y}+z_{i,j}^{yT}u^{y}_{i}+\epsilon_{i,j}.$
We take the errors $\epsilon_{i,j}$ to be normally distributed. This model implies that $y_{i,j}\sim\mathcal{N}(x_{i,j}^{yT}\beta^{y}+z_{i,j}^{yT}u^{y}_{i},\omega^{2})$, where the standard deviation $\omega$ in the normal distribution $\mathcal{N}$ is constant for all observations (homoscedasticity). The MELSM relaxes this assumption by specifying a model for the scale, which can depend on random effects, as well as covariates, i.e.:
$\omega_{i,j}=\exp(x_{i,j}^{\omega T}\beta^{\omega}+z_{i,j}^{\omega T}u^{\omega}_{i})$
where $\beta^{\omega}$ specifies the contribution of the different covariates to the variance. We model the logarithm of the standard deviation to ensure non-negativity. As both location and scale present fixed effects (denoted by $\beta$) and random effects (denoted by $u_{i}$), we distinguish the different parameters using superscripts $\omega$ and $y$. Note that, for the normal distribution, the location and scale correspond to the traditional mean and standard deviation. In the remainder, we use the terms location and scale to distinguish the submodels’ parameters.
The random effect vectors $u^{y}_{i}$ and $u^{\omega}_{i}$ are to be interpreted as subject-specific deviations from the population mean and standard deviation, respectively. For ease of notation, random intercepts $u^{y}_{0,i}$ and $u^{\omega}_{0,i}$ are implicitly included by setting the first column of the corresponding covariate matrices to a vector of 1s, while random slopes $u^{y}_{1,i}$ and $u^{\omega}_{1,i}$ are obtained by setting the column to the vector of observation times. The random effects for the location and scale models are assumed to be normally distributed with zero mean, and a correlation between random effects can be estimated. For example, in the simple setting with just a random intercept for both the location and scale models, we can specify the correlation $\rho$ to capture the relationship between the random intercept for the mean and the random intercept for the variance:
$\begin{pmatrix}u_{i}^{y}\\ u_{i}^{w}\end{pmatrix}\sim\mathcal{N}\begin{bmatrix}\begin{pmatrix}0\\ 0\end{pmatrix}\!\!,\begin{pmatrix}\sigma_{y}^{2}&\rho\sigma_{y}\sigma_{\omega}\\ \rho\sigma_{y}\sigma_{\omega}&\sigma_{\omega}^{2}\end{pmatrix}\end{bmatrix}$ (1)
@@ 3.1 Data generation
We generate simulation data based on the Primary Biliary Cholangitis (PBC) dataset, publicly available in the R package survival and commonly used as a test for longitudinal modelling. The dataset results from a clinical trial at Mayo Clinic studying the impact of D-penicillamine on PBC, an autoimmune condition affecting the liver. From the dataset, we extract four covariates at baseline: age, albumin, triglycerides (trig), and platelets.
Using these covariates, we generate a synthetic population of $N$ individuals by drawing from a normal distribution with the covariance defined by the previously extracted and standardised covariates at baseline. Note that $N$ is a parameter of the simulation and not the observed PBC population size, as one of our aims is to study how an increasing number of patients impacts our conclusions. Formally, we draw $x_{i,0}$ – where $i$ is the index associated with the patient and 0, the baseline encounter – from a multivariate normal distribution $\text{MVN}(0,\Sigma)$ with the entries of the covariance matrix $\Sigma$ given in Table 1.
@@
where $\mathcal{U}(a,b)$ denotes a uniform distribution with lower and upper bounds $a$ and $b$, respectively, and $M$ is a parameter of the simulation study representing the average number of encounters per subject across the population. This encounter process reflects a regular schedule uninformative of the outcome of interest. For each encounter, we compute the age at this encounter as follows:
@@
and we use this covariate as the only time-varying variable while all other covariates remain constant over time. Unless otherwise specified, we consider a total of $N=200$ patients with an average number of observations $M=15$. Further, we simulate random intercepts for all individuals defined as in Equation 1. We set $\rho=0$, $\sigma_{y}=1$ and $\sigma_{\omega}=0.5$ to induce non-negligible variability in the simulated outcomes.
@@ 3.2 Estimands of interest
Our simulations explore the recovery of the fixed and random effects for both scale and location models. To this end, we report the box plots and the associated coverage (for 95% credible intervals) for the estimated coefficients of interest across 500 repetitions of the simulation described in Section 3.1. Formally, the coverage of the parameter $\theta$ (in percent) is defined as follows:
@@
in which $\mathbbm{1}(c)$ denotes the indicator function, which equals 1 when the condition $c$ is satisfied and 0 otherwise; and $L$ and $U$ are the lower and upper bounds of the estimated $\hat{\theta}$ at the confidence level $\alpha$ computed as the quantiles of the posterior samples, meaning that for accurately estimated parameters and standard errors the coverage probability should be $\alpha$. Additional metrics (estimated credible interval width, the standard deviation of the estimates, and average estimated standard errors across the different replications) are presented in Appendix B to further support the insights of the conclusions presented below.
@@ 3.3 Software
All experiments are implemented in R with code made available on GitHub [^2] https://github.com/Jeanselme/IndividualVariability. We fit the MELSM models using the brms package [burkner2017brms], with default priors, two chains, and a burn-in of 1000 iterations for a total of 2000 iterations.
@@ 4 Results
This section describes common practices when using mixed-effect models and their impact on parameter’ estimates. For each model, we report the formula in pseudo-R code for reproducibility.
@@ 4.1 Practice 1: Ignoring heteroscedasticity
By relying on LMM, practitioners typically assume homoscedasticity, thereby ignoring potential subject-specific differences in variance. In this first setting, we compare an LMM, assuming a constant scale, with the properly specified MELSM model. The two models share the same (correct) specification for the location, with only the scale differing:
@@
In this first set of simulations, we investigate the effect of varying the number of patients and observations per patient, i.e., $N\in\{100,300,500,1000\}$ and $M\in\{5,10,20\}$. These values were chosen to approximate different data amounts and frequencies as common in real longitudinal data settings. Figure 1 presents the estimated fixed effects associated with age in the location and scale models under varying numbers of individuals $N$ across 100 simulations. Similarly, Figure 2 presents the same results under varying average numbers of observations per patient. In each figure, the red vertical line indicates the true value of the estimand.
@@
The coverage in the LMM is well below the nominal level of 95%, indicating that the standard errors are underestimated. Consequently, the misspecification of the scale has a direct impact on the inference for the location parameters.
Larger numbers of individuals and encounters reduce error.
Both models benefit from a larger number of individuals and encounters, as shown by reduced variance in the location estimates and reduced variance for MELSM scale estimates.
@@ 4.2 Practice 2: Misspecifying location and scale in MELSM
Figure 3: Estimated age effect for location (left panel) and scale (right panel). For each model, coverage is reported as a percentage next to the corresponding boxplot.
Figure 4: Estimated standard deviation of random intercepts for location (left panel) and scale (right panel). For each model, coverage is reported as a percentage next to the corresponding boxplot.
Figure 3 presents the fixed effect estimates associated with the age covariate, while Figure 4 describes the estimated standard deviation of the random intercept across the simulations. On the right, we display the coverage in percent over the different simulations. For consistency, we display all method specifications in these figures. Note that some do not estimate all quantities, hence the absence of some estimates.
@@
In all figures, the correctly specified model (Correct) and the one considering all covariates (All) result in similar estimates. This indicates that, with sufficient data, all models that include the correct covariates as a special case will yield a correct inference. This is also confirmed by the other metrics in Appendix B. Note, however, that considering all covariates results in an increased number of parameters and associated computational time; the fitting procedure for the method considering all covariates is more than 50% slower than the one for the correctly specified model.
@@
The age effect in the location model is well estimated across all models, although the variance of the estimates increases as the scale model departs from the true one (as observed with LMM). The largest variance is observed for No $u^{\omega}$ + Mis. $\omega$, where both a relevant covariate and the random intercept are missing in the scale model. The same model also shows, on average, lower estimated standard deviations for the location random intercept than the true value.
@@
In the misspecified location model Mis. $y$, the scale fixed effect estimate and random intercept are biased, despite presenting the right scale specification. As the location parameterisation does not capture the full variation of the data, the residual variance differs from the true value, and therefore, all scale estimates are biased (Appendix A provides a theoretical intuition of why this occurs).
@@
The estimated standard deviation of location random intercept can incur bias when the scale model is incorrectly specified (and similarly when a LMM is wrongly used in place of a MELSM). The standard deviation for the random effect of the scale model is biased in all considered scenarios unless the fixed effect structure for both location and scale is correctly specified.